# Quark-versus-gluon tagging in CMS Open Data with CWoLa and TopicFlow

### v1

@@ Appendix B Comparison with prior work
Here we present validation of our workflow by comparing both the data and simulation datasets yielded by our pipeline to those used in Refs. [54], made accessible through the EnergyFlow package [106] and various Zenodo records [107, 108, 109, 110, 111, 112, 113, 114, 115].
@@
The authors of Refs. [54] base their analysis on $R=0.5$ anti-$k_{t}$ jets sourced from the 2011 Jet Primary Dataset and the corresponding QCD simulation samples from CMS. Their focus is on events firing the single-jet Jet300 high-level trigger, which is unprescaled. To facilitate a comparison of our processing of the data, we isolate events from this trigger in our raw data – that is, the data output by our custom EDAnalyzer, discussed above – and apply a consistent set of baseline selection criteria to both datasets. Specifically, we require jets with $|\eta|<1.9$ and $p_{T}\in[375,1000]~{}\mathrm{GeV}$, and that pass medium JQC.
@@
In Fig. 10 we show a comparison of jet kinematics, the jet multiplicity and JEC values for data from the CMS Jet Primary Dataset, from both our data pipeline and that of the MIT group. Following Ref. [54], we keep only the hardest two jets in each event for this comparison. We see that there is good agreement in all of the histograms shown. Additionally, in Fig. 11a we show the distribution of the number of primary vertices for the same datasets. Here we note a discrepancy between the distributions, which we understand as coming about from the additional requirements we place on the vertex collection sourced from CMS, described in detail in Sec. A.2. For each individual event, our dataset consistently exhibits an equal or lower count of primary vertices compared to the MIT data. The ratio of the NPV distributions in data and simulation is used to mitigate pileup effects through a reweighting of MC histograms. We have checked that the NPV weights derived from the MIT data — or equivalently, derived from our pipeline without the additional requirements on the vertex collection — lead to negligible differences in the distributions of kinematic and jet-substructure variables.
@@
Figure 10: The figure shows a comparison between data sourced from our data pipeline and those from the MIT group [54].
Figure 11: The figure shows a comparison of the distribution of the number of primary vertices in (a) real data and (b) simulation between our data pipeline and the data published by the MIT group [54].
Figure 12: The figure shows a comparison between synthetic data sourced from our data pipeline and those from the MIT group [54]. For the case of the constituent multiplicity we have distinguished between quarks (green) and gluons (purple).

### v2

@@ Appendix D Comparison with prior work
Here we present validation of our workflow by comparing both the data and simulation datasets yielded by our pipeline to those used in Refs. Komiske:2019jim , made accessible through the EnergyFlow package energyflow and various Zenodo records komiske2019a ; komiske2019b ; komiske2019c ; komiske2019d ; komiske2019e ; komiske2019f ; komiske2019g ; komiske2019h ; komiske2019i .
@@
The authors of Refs. Komiske:2019jim base their analysis on $R=0.5$ anti-$k_{t}$ jets sourced from the 2011 Jet Primary Dataset and the corresponding QCD simulation samples from CMS. Their focus is on events firing the single-jet Jet300 high-level trigger, which is unprescaled. To facilitate a comparison of our processing of the data, we isolate events from this trigger in our raw data – that is, the data output by our custom EDAnalyzer, discussed above – and apply a consistent set of baseline selection criteria to both datasets. Specifically, we require jets with $|\eta|<1.9$ and $p_{T}\in[375,1000]~\mathrm{GeV}$, and that pass medium JQC.
@@
In Fig. 12 we show a comparison of jet kinematics, the jet multiplicity and JEC values for data from the CMS Jet Primary Dataset, from both our data pipeline and that of the MIT group. Following Ref. Komiske:2019jim , we keep only the hardest two jets in each event for this comparison. We see that there is good agreement in all of the histograms shown. Additionally, in Fig. 13a we show the distribution of the number of primary vertices for the same datasets. Here we note a discrepancy between the distributions, which we understand as coming about from the additional requirements we place on the vertex collection sourced from CMS, described in detail in Sec. A.2. For each individual event, our dataset consistently exhibits an equal or lower count of primary vertices compared to the MIT data. The ratio of the NPV distributions in data and simulation is used to mitigate pileup effects through a reweighting of MC histograms. We have checked that the NPV weights derived from the MIT data — or equivalently, derived from our pipeline without the additional requirements on the vertex collection — lead to negligible differences in the distributions of kinematic and jet-substructure variables.
@@
Figure 12: The figure shows a comparison between data sourced from our data pipeline and those from the MIT group Komiske:2019jim .
Figure 13: The figure shows a comparison of the distribution of the number of primary vertices in (a) real data and (b) simulation between our data pipeline and the data published by the MIT group Komiske:2019jim .
Figure 14: The figure shows a comparison between synthetic data sourced from our data pipeline and those from the MIT group Komiske:2019jim . For the case of the constituent multiplicity we have distinguished between quarks (green) and gluons (purple).